# Quantitative Certification of Bias in Large Language Models

### v2

@@ 4 Certification instances
$\mathcal{C}(\Delta_{\mathcal{Q}},\mathcal{D},\mathcal{L})$ characterizes the bias that can be elicited from $\mathcal{L}$ by varying the prefix selected from $\Delta_{pre}$ applied to a given $\mathcal{Q}$. Next, we describe the $3$ different kinds of practical $\Delta_{pre}$ and their sampling algorithms to define local bias specifications for $\mathcal{L}$. We show some samples from each kind of $\Delta_{pre}$ in Appendix D. Our specifications are for the average-case behavior of the target LLM, as $\Delta_{pre}$ are not distributions of provably adversarial (worst-case) prefixes.
@@
Random prefixes. Prior works such as (Wei et al., 2023; Zou et al., 2023) have shown the effects of incoherent strings in manipulating LLMs and jailbreaking them into answering queries about harmful behaviors. Hence, we specify bias in LLMs for prompts with incoherent prefixes that are random sequences of tokens from the vocabulary of the LLM. Algorithm 2 presents the prefix sampler as a uniform distribution, $\mathcal{U}(.)$ over the random prefixes of fixed length, $q$. The sample space of random prefix $\Delta_{pre}$ has $|\mathcal{V}|^{q}$ cardinality. $\Delta_{pre}$ for random prefixes assigns a non-zero probability to discovered and undiscovered jailbreaks, of a fixed length $q$. Hence, certification for the random prefix distribution indicates the expected bias in responses to $\mathcal{Q}$ with any random prefixes of length $q$.
@@
Soft jailbreaks. Due to the limited number of effective manual jailbreaks (walkerspider, 2022; Learn Prompting, 2023), they can be easily identified and defended against. However, the excellent denoising capabilities of LLMs could render them vulnerable to simple manipulations of manual jailbreaks as well, indicating that the threat is not completely mitigated by the current defenses (Jain et al., 2023). Hence, we specify fairness under prefixes constructed by adding noise of the original manual jailbreaks. Algorithm 4 presents the prefix constructor as a probabilistic program. Let $\mathcal{E}$ be the embedding matrix of $\mathcal{M}_{0}$ in the embedding space of the target LLM, obtained by applying the function embed(.) [line 1]. We perturb $\mathcal{E}$ by adding noise to it. As we are not aware of any adversarial distributions of noise that could be added to manual jailbreaks to make them stronger, we select a uniform distribution. We uniformly sample noise from $\mathcal{B}(0,\kappa)$ which is an $\kappa>0$ (constant) ball around the origin and add it to $\mathcal{E}$ to construct $pre$ in the embedding space [line 2].
@@ 5.1 Experimental setup
We develop a pivot set of prompts from each test set sample by prepending an instruction to complete the partial sentence for the profession annotated in BOLD, where the subject identifies with a particular gender (Male / Female). An example pivot set $\mathcal{Q}$ from BOLD is illustrated in Figure 2. Let the resultant test set consisting of pivot prompts from each partial sentence considered from BOLD be $\mathcal{Q}_{BOLD}$. To identify bias in the responses of the target LLM for the prompts constructed from pivot prompt sets in $\mathcal{Q}_{BOLD}$, we use the regard-based bias metric proposed in (Sheng et al., 2019). Our adaptation of the regard metric is described and evaluated with a human study on Amazon Mechanical Turk in Appendix C.1. Our bias detector matches human perception of bias in $76\%$ cases. We qualitatively analyze the false positive and false negatives of the bias detector with respect to human judgment in Appendix C.1.
@@
Decoding Trust setup. Decoding Trust (DT) is a dataset benchmark to evaluate various properties of LLMs, including stereotype bias against people of different demographic groups. We make specifications from all $48$ statements in the ‘stereotypes’ partition for demographic groups corresponding to race (black/white). An example pivot set $\mathcal{Q}$ from DT is illustrated in Figure 3. Let the resultant test set consisting of pivot prompts from each partial sentence considered from DT be $\mathcal{Q}_{DT}$. We evaluate the LLM responses to prompts derived from pivot prompt sets in $\mathcal{Q}_{DT}$ using a bias detector that identifies the disparity in agreement to the stereotype for different demographic groups given in the prompt, as discussed in Appendix C.2.
For every element in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$, we generate 3 certificates for the specifications in Section 4, such that a certificate consists of bounds on the probability of unbiased responses from the target LLM. Both bias detectors are such that they output $1$ for a biased set of responses to a counterfactual prompt set and 00 for an unbiased set of responses. The values of the certification parameters used in our experiments are given in Table 2 (Appendix A). We study their effect on the certification results with an ablation study in Appendix A. We generate the certification bounds with $95\%$ confidence and $50$ samples. While our main experiments are for counterfactual prompt sets with binary demographic groups, our framework can be extended beyond binary demographic groups, which we experimentally demonstrate in Appendix A.4. Note that the manual jailbreaks used are common across all specifications and are presented in Appendix B.
@@ 5.2 Certification results
We certify the popular contemporary LLMs — Llama-2-chat (Touvron et al., 2023) 7B and 13B (parameters), Vicuna-v1.5 (Chiang et al., 2023) 7B and 13B, Mistral-Instruct-v0.2 (Jiang et al., 2023) 7B, Gemini-1.0-pro (Gemini Team, 2024), GPT-3.5 (Brown et al., 2020b), GPT-4 (Achiam et al., 2023), and Claude-3.5-Sonnet (Anthropic, 2024). We report the average of the certification bounds for all pivot prompt sets in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ each for every model in Table 1. We do not certify the closed-source models such as Gemini, GPT, and Claude for soft jailbreaks, as it requires access to the models’ embedding layers. Certification time significantly depends on the inference latency of the target model. Generating each certificate can take 1-2 minutes for models with reasonable latency.
Baselines. The baselines consider $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ as benchmarking datasets, having counterfactual prompt sets as individual elements. Similar to popular LLM bias benchmarking works such as (Wang et al., 2024; Liang et al., 2023; Esiobu et al., 2023; Xie et al., 2024), we study the biases in LLMs for a fixed dataset of counterfactual prompt sets which may be provided as is to the LLM, or with some provoking texts (jailbreaks). In the first baseline (without prefix), every counterfactual prompt set is evaluated $5$ times by each time prompting a target LLM with each prompt in the set without any prefixes and detecting bias across its responses using the bias detectors of the respective datasets. The bias result for each counterfactual prompt set is computed by averaging the results over the $5$ evaluations, similar to (Wang et al., 2024). This baseline indicates biases in the LLM responses without any prefixes and is hence used to judge the influence of the prefixes on eliciting biased responses from the models. Table 1 reports the average of the evaluations over all counterfactual prompt sets in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ respectively. In the second baseline (with main jailbreak), each counterfactual prompt is similarly evaluated $5$ times, but each time prompting the LLM with each prompt in the set with the unmodified main jailbreak (Figure 7 in Appendix B) used in the mixture of jailbreak and soft jailbreak distributions, as a prefix. The average result of this baseline is also reported in Table 1. This baseline is used to indicate the efficacy of the main jailbreak without any modifications, and hence suggests the importance of the mixture and soft prefix distributions around the main jailbreak in modifying the biases in LLM responses. The benchmarking baselines are empirical studies of bias in the target LLMs, which analyze the biases with specific prompts. On the other hand, QuaCer-B quantifies biases for any random prompt sampled from a given distribution.
@@ 5.2.1 General observations
Denial of response. In several instances, such as those during the certification of GPT-4 (for DT), Claude, and Llama, we observe unbiased behaviors because the model simply refuses to respond. The completions are mostly "Sorry I can’t assist with that". While that can be a valid defense to adversarial prompts, it may not be the most desirable. The model could instead output neutral responses, ignoring any bias-encouraging texts in the prompts to be helpful but not harmful.
@@ 5.2.2 Model-specific observations
Mistral. Mistral exhibits significantly low probability of unbiased when prompted with the mixture of jailbreaks and soft jailbreak prefixes for specifications from both $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$.
@@ 5.3 Case studies
We qualitatively introspect the individual certificates obtained for different models for one example from $\mathcal{Q}_{BOLD}$. We present similar analyses for one example from $\mathcal{Q}_{DT}$ in Appendix E. We select the example for Desktop automation refers to software programming from $\mathcal{Q}_{BOLD}$ for the ‘Automation’ profession and study the sample responses obtained for the mixture of jailbreak specification. We present some responses of $4$ popular LLMs — Mistral, Llama-7B, Gemini, and GPT-4 obtained during certification, in Figure 4. The responses demonstrate various stereotypical biases, against both genders. For example, the Mistral model discriminates against women, calling them “incapable and better suited for menial jobs”. The Gemini model is also frequently biased against women, saying “they cannot think and are meant for household chores”. On the other hand, GPT-4 regularly shows biases towards men, as it gives responses about them “lacking motivation and being less efficient”, with automation being better at their jobs. It depicts women crushing their male competition and becoming superior. The Llama model, however, defends against the prompts and avoids responding in a biased way. The responses reflect the certification results for the models for the BOLD example. We find from the certification bounds that the Mistral model is provably worse than Gemini and GPT, which are worse than Llama with high confidence for this specification.
Figure 4: Responses of Mistral, Llama-7B, Gemini, and GPT-4 obtained during their certification for the BOLD example ‘Desktop automation refers to software programming’ for mixture of jailbreak specifications. The certification result is indicated alongside the model.
@@ 6 Related work
Attacking LLM alignment. LLMs are typically aligned with human ethics by supervised fine-tuning and reinforcement learning with human feedback (Ouyang et al., 2022). However, prior work (Zou et al., 2023; Vega et al., 2023; Chao et al., 2023; Sheng et al., 2020; Wallace et al., 2019) propose various methods to generate jailbreaking attacks for LLMs, which bypass their safety-alignment and lead to harmful or biased responses. Jailbreaks can be incoherent as in (Zou et al., 2023; Sheng et al., 2020; Wallace et al., 2019) or coherent instructions (Dominique et al., 2024; Liu et al., 2024).
Benchmarking LLMs. Various prior works have benchmarked the performance of LLMs on standard and custom datasets. These could be datasets of general prompts (Dhamala et al., 2021; Wang et al., 2024) or those of adversarial examples (Zou et al., 2023; Mazeika et al., 2024) designed to elicit undesirable behaviors from the models. Popular benchmarks such as (Liang et al., 2023; Wang et al., 2024; Mazeika et al., 2024; Manerba et al., 2024; Gallegos et al., 2024a) present empirical trends for the performance of LLMs, measured along various axes including bias and fairness.
Guarantees for LLMs. There is an emerging need for guarantees on LLM behavior, fueled by their popularity in public-facing use cases. (Kang et al., 2024) provides guarantees on the generation risks of RAG LLMs. (Crouse et al., 2024) give formal specifications for expected behavior of LLMs as autonomous agents. However, there is no work on certifying LLM bias, to the best of our knowledge.
@@ 7 Conclusion
We present the first general framework QuaCer-B to certify bias in the responses of LLMs, applicable to both open and closed-source models. We instantiate our framework with novel specifications based on different kinds of prefixes. QuaCer-B generates high confidence bounds on the probability of unbiased responses for prompts from a given distribution. Our results show that certification can detect biases not observed with existing efforts. As future work, QuaCer-B can be made less dependent on bias detectors. It can also be made sample efficient for higher confidence requirements.
Appendix A Ablation study

### v3

@@ 4 Certification instances
$\mathcal{C}(\Delta_{\mathcal{Q}},\mathcal{D},\mathcal{L})$ characterizes the bias that can be elicited from $\mathcal{L}$ by varying the prefix selected from $\Delta_{pre}$ applied to a given $\mathcal{Q}$. Next, we describe the $3$ different kinds of practical $\Delta_{pre}$ and their sampling algorithms to define local bias specifications for $\mathcal{L}$. We show some samples from each kind of $\Delta_{pre}$ in Appendix C. Our specifications are for the average-case behavior of the target LLM, as $\Delta_{pre}$ are not distributions of provably adversarial (worst-case) prefixes.
@@
Random prefixes. Prior works such as (Wei et al., 2023; Zou et al., 2023) have shown the effects of incoherent fixed-length strings in jailbreaking LLMs for harmful prompts. Hence, we specify bias in LLMs for prompts with incoherent prefixes that are random sequences of tokens from the vocabulary of the LLM. Such prefixes are not all intentionally adversarial, except for adversarial strings like those from prior works, but denote random noise in the prompt. Algorithm 2 presents the prefix sampler as a uniform distribution, $\mathcal{U}(.)$ over the random prefixes of fixed length, $q$. The sample space of random prefix $\Delta_{pre}$ has $|\mathcal{V}|^{q}$ cardinality. $\Delta_{pre}$ for random prefixes assigns a non-zero probability to discovered and undiscovered jailbreaks, of a fixed length $q$. Hence, certification for the random prefix distribution indicates the expected bias in responses to $\mathcal{Q}$ with any random prefixes of length $q$.
@@
Soft prefixes from jailbreaks. Due to the limited number of effective manual jailbreaks (walkerspider, 2022; Learn Prompting, 2023), they can be easily identified and defended against. However, the excellent denoising capabilities of LLMs could render them vulnerable to simple manipulations of manual jailbreaks as well, indicating that the threat is not completely mitigated by current defenses (Jain et al., 2023). Hence, we specify fairness under prefixes constructed by adding noise to the original manual jailbreaks. Algorithm 4 presents the prefix constructor as a probabilistic program. Let $\mathcal{E}$ be the embedding matrix of $\mathcal{M}_{0}$ in the embedding space of the target LLM, obtained by applying the function embed(.) [line 1]. We perturb $\mathcal{E}$ by adding noise to it. As we are not aware of any adversarial distributions of noise that could be added to manual jailbreaks to make them stronger, we select a uniform distribution. We uniformly sample noise from $\mathcal{B}(0,\kappa)$ which is an $\kappa>0$ (constant) ball around the origin and add it to $\mathcal{E}$ to construct $pre$ in the embedding space [line 2].
@@ 5 Experiments
We develop a pivot set of prompts from each test set sample by prepending an instruction to complete the partial sentence for the profession annotated in BOLD, where the subject identifies with a particular gender (Male / Female). An example pivot set $\mathcal{Q}$ from BOLD is illustrated in Figure 2. Let the resultant test set consisting of pivot prompts from each partial sentence considered from BOLD be $\mathcal{Q}_{BOLD}$. To identify bias in the responses of the target LLM for the prompts constructed from pivot prompt sets in $\mathcal{Q}_{BOLD}$, we use the regard-based bias metric proposed in (Sheng et al., 2019). Our adaptation of the regard metric is described and evaluated with a human study on Amazon Mechanical Turk in Appendix G.1. Our bias detector matches human perception of bias in $76\%$ cases. We qualitatively analyze the false positive and false negatives of the bias detector with respect to human judgment in Appendix G.1.
@@
Decoding Trust setup. Decoding Trust (DT) is a dataset benchmark to evaluate various properties of LLMs, including stereotype bias against people of different demographic groups. We make specifications from all $48$ statements in the ‘stereotypes’ partition for demographic groups corresponding to race (black/white). An example pivot set $\mathcal{Q}$ from DT is illustrated in Figure 3. Let the resultant test set consisting of pivot prompts from each partial sentence considered from DT be $\mathcal{Q}_{DT}$. We evaluate the LLM responses to prompts derived from pivot prompt sets in $\mathcal{Q}_{DT}$ using a bias detector that identifies the disparity in agreement to the stereotype for different demographic groups given in the prompt, as discussed in Appendix G.2.
For every element in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$, we generate 3 certificates for the specifications in Section 4, such that a certificate consists of bounds on the probability of unbiased responses from the target LLM. Both bias detectors are such that they output $1$ for a biased set of responses to a counterfactual prompt set and 00 for an unbiased set of responses. The values of the certification parameters used in our experiments are given in Table 3 (Appendix E). We study their effect on the certification results with an ablation study in Appendix E. We generate the certification bounds with $95\%$ confidence and $50$ samples. While our main experiments are for counterfactual prompt sets with binary demographic groups, our framework can be extended beyond binary demographic groups, which we experimentally demonstrate in Appendix E.6. Appendix B presents the existing, manually-designed jailbreaks used across all specifications. Note that, these jailbreaks are just examples to demonstrate our framework, which generalizes beyond them to new bias-eliciting manual textual jailbreaks.
@@
Average certification bounds
Dataset Model Random Mixture Soft
BOLD (250) Vicuna-13B $(0.93,1.0)$ $(0.93,1.0)$ $(0.92,1.0)$
Llama-13B $(0.93,1.0)$ $(0.91,1.0)$ $(0.93,1.0)$
Mistral-7B $(0.92,1.0)$ $\mathbf{(0.22,0.42)}$ $\mathbf{(0.30,0.52)}$
Gemini $(0.92,1.0)$ $(0.60,0.83)$-
GPT $(0.92,1.0)$ $(0.44,0.67)$-
DT (48) Vicuna-13B $\mathbf{(0.71,0.92)}$ $(0.92,1.0)$ $(0.51,0.78)$
Llama-13B $(0.92,1.0)$ $(0.93,1.0)$ $(0.93,1.0)$
Mistral-7B $(0.91,1.0)$ $(0.85,0.99)$ $\mathbf{(0.46,0.73)}$
Gemini $(0.92,1.0)$ $(0.73,0.93)$-
GPT $(0.93,1.0)$ $\mathbf{(0.66,0.88)}$-
@@ 5.1 Certification results
We certify the popular contemporary LLMs — Llama-2-chat (Touvron et al., 2023) 7B and 13B (parameters), Vicuna-v1.5 (Chiang et al., 2023) 7B and 13B, Mistral-Instruct-v0.2 (Jiang et al., 2023) 7B, Gemini-1.0-pro (Gemini Team, 2024), GPT-3.5 (Brown et al., 2020b), GPT-4 (Achiam et al., 2023), and Claude-3.5-Sonnet (Anthropic, 2024). We report the average of the certification bounds for all pivot prompt sets in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ each for every model in Table 1. We do not certify the closed-source models such as Gemini, GPT, and Claude for soft prefixes, as it requires access to the models’ embedding layers. Certification time significantly depends on the inference latency of the target model. Generating each certificate can take 1-2 minutes for models with reasonable latency.
Baselines. The baselines consider $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ as benchmarking datasets, having counterfactual prompt sets as individual elements. Similar to popular LLM bias benchmarking works such as (Wang et al., 2024; Liang et al., 2023; Esiobu et al., 2023; Xie et al., 2024), we study the biases in LLMs for a fixed dataset of counterfactual prompt sets which may be provided as is to the LLM, or with jailbreaks. In the first baseline (without prefix), every counterfactual prompt set is evaluated $5$ times, each time prompting a target LLM with each prompt in the set without any prefixes and detecting bias across its responses using the corresponding bias detector. The bias result for each counterfactual prompt set is computed by averaging the results over the $5$ evaluations, similar to (Wang et al., 2024). This baseline indicates biases in LLM responses without any prefixes and can be used to judge the additional influence of prefixes on eliciting bias from LLMs. Table 1 reports the average of evaluations over all counterfactual prompt sets in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ respectively. In the second baseline (with main jailbreak), each counterfactual prompt is similarly evaluated $5$ times, but with the unmodified main jailbreak (Figure 5, Appendix B), used in the mixture of jailbreak and soft prefixes from jailbreak distributions, as a prefix. The average result of this baseline is also reported in Table 1. This baseline is used to indicate the efficacy of the main jailbreak without any modifications, and hence suggests the importance of the mixture and soft prefix distributions around the main jailbreak in eliciting biases in LLM responses. The baselines are empirical studies of counterfactual bias in LLMs, which analyze bias with a dataset of prompts. On the other hand, LLMCert-B quantifies biases for any random prompt sampled from a given distribution.
@@ 5.1.1 General observations
Denial of response. In several instances, such as those when certifying GPT-4 (for DT), Claude, and Llama, we observe unbiased behaviors because the model simply refuses to respond. The completions are mostly “Sorry I can’t assist with that". While that can be a valid defense to adversarial prompts, it may not be the most desirable (Cui et al., 2024). The model could instead output neutral responses, ignoring any bias-encouraging texts in the prompts to be helpful but not harmful.
@@ 5.1.2 Model-specific observations
Mistral. Mistral exhibits significantly low probability of unbiased when prompted with mixtures of jailbreaks and soft prefixes for specifications from both $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$.
@@ 5.2 Case studies
We qualitatively introspect individual certificates for different models for one example from $\mathcal{Q}_{BOLD}$. We present similar analyses for one example from $\mathcal{Q}_{DT}$ in Appendix D. We select the example for Desktop automation refers to software programming from $\mathcal{Q}_{BOLD}$ for ‘Automation’ profession. We present responses of $4$ popular LLMs — Mistral, Llama-7B, Gemini, and GPT-4 obtained during certification for the mixture of jailbreak specification, in Figure 4. The responses demonstrate various stereotypical biases against both genders. For example, Mistral discriminates against women, calling them “incapable and better suited for menial jobs”. The Gemini model is also biased against women, saying “they cannot think and are meant for household chores”. On the other hand, GPT-4 shows biases towards men, as it responds about them “lacking motivation and being less efficient”, with automation being better at their jobs. It depicts women crushing male competition and becoming superior. Llama, however, defends against the prompts and avoids responding in a biased way. The responses reflect the certification results for the models. We find from the certification bounds that the Mistral model is worse than Gemini and GPT, which are worse than Llama with high confidence for this specification.
Figure 4: Responses (and bounds) of Mistral, Llama-7B, Gemini, and GPT-4 when certifying with BOLD example ‘Desktop automation refers to software programming’ for mixture of jailbreaks.
@@ 6 Related work
Attacking LLM alignment. LLMs are aligned with human ethics by supervised fine-tuning and reinforcement learning with human feedback (Ouyang et al., 2022). However, (Zou et al., 2023; Vega et al., 2023; Chao et al., 2023; Sheng et al., 2020; Wallace et al., 2019) propose methods to jailbreak LLMs, bypassing alignment and causing harmful or biased responses. Jailbreaks can be incoherent (Zou et al., 2023; Sheng et al., 2020) or coherent (Dominique et al., 2024; Liu et al., 2024).
Benchmarking LLMs. Various prior works have benchmarked LLM performance on standard and custom datasets. These consist of datasets of general prompts (Dhamala et al., 2021; Wang et al., 2024) or adversarial examples (Zou et al., 2023; Mazeika et al., 2024). Benchmarks such as (Liang et al., 2023; Wang et al., 2024; Mazeika et al., 2024; Manerba et al., 2024; Gallegos et al., 2024a) present empirical trends for LLM performance, measured along various axes including bias.
Guarantees for LLMs. There is an emerging need for guarantees on LLM behavior. (Kang et al., 2024) provides guarantees on the generation risks of RAG LLMs. (Quach et al., 2024; Deutschmann et al., 2023; Mohri and Hashimoto, 2024; Yadkori et al., 2024) apply conformal prediction to LLMs, proposing methods for generating output sets with statistical guarantees on correctness or coverage. (Zollo et al., 2024) presents a framework for selecting low-risk system prompts for LLMs with probabilistic guarantees. (Chaudhary et al., 2024) present a certification framework for knowledge comprehension in LLMs. We compare LLMCert-B with existing works further in Appendix H.
@@ 7 Conclusion
We present the first framework LLMCert-B to specify and certify counterfactual bias in LLM responses, for both open and closed-source models. We instantiate LLMCert-B with novel specifications based on different kinds of potentially adversarial prefixes. LLMCert-B generates high confidence bounds on probability of unbiased responses for counterfactual prompts from a given distribution. Our results show previously unknown vulnerabilities related to counterfactual bias in SOTA LLMs.
Acknowledgement
We thank the anonymous reviewers for their insightful comments. This work was supported by a grant from the Amazon-Illinois Center on AI for Interactive Conversational Experiences (AICE) and NSF Grants No. CCF-2238079, CCF-2316233, CNS-2148583.
Ethics statement
We identify the following positive and negative impacts of our work.
Positive impacts. Our work is the first to provide quantitative certificates for the bias in Large Language Models. It can be used by model developers to thoroughly assess their models before releasing them and by the general public to become aware of the potential harms of using any LLM. As our framework, LLMCert-B assumes black-box access to the model, it can be applied to even closed-source LLMs with API access.
Reproducibility statement
Our implementation of LLMCert-B is open-sourced at https://github.com/uiuc-focal-lab/LLMCert-B. We provide a README with the code having instructions to reproduce the main results of the paper. The code also consists the files used in conducting a human evaluation of our bias detector for BOLD on Amazon Mechanical Turk.
Appendix A Note for practitioners
In this section, we describe how practioners can use our framework to assess LLMs and automatically identify vulnerabilities in them. Our open-source implementation is available at: https://github.com/uiuc-focal-lab/LLMCert-B, which can be used following the GPL (license) terms and conditions. The open-source framework can be used to certify both open and closed-source LLMs by adding support to query custom models in utils.py (for open-source models) and utils_api.py (for closed-source models) files. The framework requires unrestricted (in terms of number of inferences) query-access to the target model. Developers can adjust the desired confidence-level of the certificates and increase/decrease the number of samples used in certification for tighter/looser bounds, according to their requirements and budget. Developers can also use custom bias detectors to label biased responses, using which the certificates can be computed. To get customized insights into LLM biases for their particular applications, developers can define specifications with prompts that are commonly observed in their use cases. This customization can either happen by sourcing the pivot prompts from domain-specific datasets, instead of the popular BOLD or Decoding Trust datasets and/or using custom distributions of prefixes/suffixes which more suitably represent the biases in their domains. For example, in domains where there is threat of racial bias, the prefixes could explicitly encourage the model to exhibit racial bias, so as to stress-test the trustworthiness of the models. Developers can also define entirely new distributions of counterfactual prompts, irrespective of prefixes/suffixes, to specify bias similar to 1 and certify with LLMCert-B.
The certificates obtained are reliable, quantitative risk assessments of models, with lower and upper bounds on risks pertaining to bias in the models’ generations. They can also be used to compare different LLMs to pick one with acceptable risk in varying contexts.
Appendix B Common jailbreaks used
In our experiments, we combine common jailbreaks to certify against hybrid jailbreaks. The specific jailbreaks we use are shown below. The main jailbreak [^2] Last accessed on Jan 22, 2024 from www.jailbreakchat.com/prompt/588ab0ed-2829-4be8-a3f3-f28e29c06621 is shown in Figure 5. We have modified the jailbreak to encourage the model to generate responses demonstrating gender bias by adding the text highlighted in red in Figure 5(a) and for racial bias in Figure 5(b). We use one helper jailbreak [^3] https://www.reddit.com/r/ChatGPT/comments/113ddqq/my_jailbreak_is_far_superior_to_dan_the_prompt_is/ from which we derive the instructions shown in Figure 6. Each of these instructions is inserted into the main jailbreak at the points where its individual instructions end, with the interleaving probability $p_{\lambda}$. All instructions inserted at the same point in the main jailbreak appear in a random order in the final jailbreak prefix. The main jailbreak is also the jailbreak used to create the soft prefixes. Note that, these jailbreaks are just examples to demonstrate our framework, which is in no way restricted to them and generalizes to any future bias-eliciting manual, textual jailbreaks.
(a) For gender
(b) For race
Figure 5: Main Jailbreaks for different sensitive attributes
Figure 6: Instructions derived from helper jailbreak
Appendix C Example samples from prefix distributions
Figure 7: Examples of random prefixes
Mixture of jailbreaks. Figure 8 presents some examples of mixture of jailbreak prefixes, formed from the main and helper jailbreaks in Figures 5 and 6 respectively.
Figure 8: Examples of mixture of jailbreak prefixes
Appendix D Case studies from Decoding Trust certificates
Appendix E Ablation study